# Can Graph Neural Networks Learn Language with Extremely Weak Text Supervision?

### v2

@@ 4 Multi-modal Prompt Learning for GNNs
As discussed in Sections 2 and 3, we introduce the text prompt $\mathbf{P}_{\theta}^{t}\in\mathbb{R}^{n_{t}\times d_{t}}$ with $n_{t}$ text prompt tokens and the graph prompt $\mathbf{P}_{\theta}^{g}\in\mathbb{R}^{n_{g}\times d}$ with $n_{g}$ graph prompt tokens. Let $\psi_{g}(\cdot,\mathbf{P}_{\theta}^{g})$ be the graph prompting function, e.g., given any graph $\mathcal{G}$, the manipulated graph $\mathcal{G}_{m}=\psi_{g}(\mathcal{G},\mathbf{P}_{\theta}^{g})$.
@@
$\omega_{t}(t,\mathbf{P}_{\theta}^{t})=[\mathbf{P}_{\theta}^{t};f_{\phi}^{t}({\rm Tokenize}(t))]_{dim=0}\in\mathbb{R}^{(len({\rm Tokenize}(t))+n_{t})\times d_{t}}$ (2)
@@
$\omega_{g}(\mathcal{G},\mathbf{P}_{\theta}^{g})=f_{\phi}^{g}(\mathcal{G}_{m})=f_{\phi}^{g}(\psi_{g}(\mathcal{G},\mathbf{P}_{\theta}^{g}))\in\mathbb{R}^{(n+n_{g})\times d_{g}}$ (3)
@@
$\mathbf{h}_{i}^{t}={\rm readout}(\omega_{t}(t_{i},\mathbf{P}_{\theta}^{t}))\in\mathbb{R}^{1\times d_{t}},\,\mathbf{h}_{i}^{\mathcal{G}}={\rm readout}(\omega_{g}(\mathcal{G}_{i},\mathbf{P}_{\theta}^{g}))\in\mathbb{R}^{1\times d_{g}}$ (4)
@@
$\mathbf{z}_{norm,i}^{\mathcal{G}}=\frac{\mathbf{h}_{i}^{\mathcal{G}}}{||\mathbf{h}_{i}^{\mathcal{G}}||_{2}}=\frac{{\rm readout}(\omega_{g}(\mathcal{G}_{i},\mathbf{P}_{\theta}^{g}))}{||{\rm readout}(\omega_{g}(\mathcal{G}_{i},\mathbf{P}_{\theta}^{g}))||_{2}},\quad\mathbf{z}_{i}^{\mathcal{G}}={\rm Proj}_{\theta}(\mathbf{z}_{norm,i}^{\mathcal{G}})$ (5)
For the text embeddings, since for limited data the set $\mathcal{T}=\{t_{i}\}_{i=1}^{C}$ may contain texts that are semantically close as discussed in Appendix B.2, we extract a subspace in the text embedding space by normalizing the embedding as follows. We further normalize the text embeddings to the unit sphere, as standard practice in NLP.
@@
Finally, we use the in-batch similarity-based contrastive loss [58] to train text prompts, graph prompts, and the projector as shown below, to adapt the pre-trained GNN representations to LLM.
@@
During inference stage, for an input graph $\mathcal{G}_{i}$ and text label candidates $\mathcal{T}=\{t_{i}\}_{i=1}^{C}$, we compute the embedding $\mathbf{z}_{i}^{\mathcal{G}}={\rm Proj}_{\theta}(\frac{{\rm readout}(\omega_{g}(\mathcal{G}_{i},\mathbf{P}_{\theta}^{g}))}{||{\rm readout}(\omega_{g}(\mathcal{G}_{i},\mathbf{P}_{\theta}^{g}))||_{2}})$ using trained $\mathbf{P}_{\theta}^{g}$ and ${\rm Proj}_{\theta}(\cdot)$. Then, we compute $\mathbf{z}^{t}_{i}$ as Equation 7 and 8. Finally, $\mathcal{G}_{i}$ will be classified to associate with text label ${\arg\max}_{1\leq i\leq C}(\mathbf{z}_{i}^{\mathcal{G}}\cdot\mathbf{z}_{i}^{t})$.
@@ 5 Experiments
GPF-plus [59] GraphCL+GCN 68.67 67.27 16.88 15.48 64.75 61.45 47.42 29.02
GraphCL+GAT 68.67 62.84 16.45 13.23 65.89 60.07 47.42 26.28
GraphCL+GT 69.33 67.87 18.12 15.56 59.66 37.37 41.71 21.35
SimGRACE+GCN 65.33 39.52 18.96 15.83 65.16 58.80 45.71 23.32
SimGRACE+GAT 69.33 66.72 18.54 12.58 63.28 53.50 42.85 21.40
SimGRACE+GT 70.00 67.31 17.91 14.69 64.83 52.97 34.13 20.13
Gprompt [60] GraphCL+GCN 73.33 66.93 17.91 8.44 61.01 60.01 1.80 0.21
GraphCL+GAT 64.67 62.63 17.08 14.18 50.56 50.55 1.80 0.22
GraphCL+GT 70.67 70.02 17.91 9.64 63.28 58.65 1.80 0.21
SimGRACE+GCN 65.33 39.52 17.29 14.48 52.70 52.68 1.80 0.21
SimGRACE+GAT 67.33 65.88 16.25 11.31 59.10 58.72 1.80 0.21
SimGRACE+GT 73.33 67.84 16.87 13.54 64.75 62.37 1.80 0.223
@@
We evaluate our Morpher and the improved graph prompt through extensive experiments. In particular, we show that, compared to state-of-the-art baseline methods, they both more effectively adapt pre-trained GNNs to the specific downstream classification task, and introducing the text modality brings Morpher additional advantages over others. We use RoBERTa [61] as the LLM encoder for Morpher in the main experiments. We also validate the performance of Morpher with ELECTRA [62] and DistilBERT [63] in section 5.6 and Appendix C.3.
Datasets. We use real-world graph datasets from PyTorch Geometric [52], including one molecular dataset MUTAG [54]; two bioinformatic datasets ENZYMES and PROTEINS [64]; one computer vision dataset MSRC_21C [65]; three citation network datasets Cora, CiteSeer and PubMed [53]. We use real-world class names as text labels. The text supervision is extremely weak, as each text label contains no more than five words. More details are summarized in Appendix B.
Pre-trained algorithms and GNN backbones. To pretrain GNNs for evaluation, we adopt GraphCL [66] and SimGRACE [67] to pre-train three widely used GNN backbones: GCN [68], GAT [69] and GraphTransformer (GT) [70]. Additionally, in Appendix C.4, we verify the effectiveness of our methods on GNNs pre-trained using GraphMAE [71] and MVGRL [72], two other representative GNN self-supervised learning algorithms. For each dataset, to pre-train GNNs, we leverage self-supervised learning methods on all the graphs without any label information.
Baselines and metrics. We compare our methods with the following baselines: (1) training a GNN from scratch supervised by few-shot data (“supervised”); (2) fine-tuning a task head together with pre-trained GNN (“fine-tune”). We allow GNNs to be tunable for “supervised” and “fine-tune”; (3) state-of-the-art graph prompting algorithms: All-in-one (“AIO”) [51], which is the only graph prompting algorithm that supports multiple tasks in node-level, edge-level and graph-level to the best of our knowledge; GPF-plus [59] which prompt on graph features and Gprompt [60] which is based on subgraph similarity. We use accuracy and weighted F1 as classification performance metrics.
@@ 5.1 Few-shot Learning
We investigate the ability of our improved graph prompt (“ImprovedAIO”) and Multimodal prompt (“Morpher”) to adapt frozen pre-trained GNNs using few-shot data. We focus on graph-level classification here and will further investigate the few-shot learning ability at other task levels in Section 5.2. Our few-shot learning setting is more challenging than existing works [51, 73] as we only allow no more than 10 labeled training and validation samples for each class. The results are shown in Table 1, where we report the average performance of 5 runs and calculate the absolute average improvement of our methods. From the results, given the same pre-trained GNN, our ImprovedAIO outperforms all the existing baseline methods except for ENZYMES F1 and MSRC_21C accuracy. Yet the performance of our ImprovedAIO on ENZYMES F1 and MSRC_21C accuracy is clearly better than those of the original AIO. Our Morpher can achieve an absolute accuracy improvement of $0.60\%$ to $7.14\%$ over the baselines across all datasets. Supervised by very limited labeled data, training a GNN from scratch is sub-optimal. Passing a GNN pre-trained on the dataset and fine-tuning it with a task head achieves sub-optimal but better results as the pre-trained GNN learns generalizable representations over the dataset through self-supervised learning. To mitigate the gap between the pre-training task and downstream tasks, AIO [51] proposes to learn graph prompts for downstream data. However, as we discussed in Section 3, when the node features are sparse vectors, the optimization would fail. Using the official implementation of AIO, we observe that the loss value tends to fluctuate, and the performance of AIO is usually even worse than supervised training. By restricting the cross-connections, our ImprovedAIO becomes more stable and constantly outperforms the fine-tuning baseline. Compared to the aforementioned methods, Morpher demonstrated superior performance due to its capability to adapt both graph and language representation spaces dynamically.
@@ 5.2 Morpher Supports Multiple-level Tasks
Table 2: Node-level, edge-level performance.
@@
Inherited from AIO, our ImprovedAIO and Morpher also support adaptation to downstream tasks at node-level and edge-level, because they can be reformulated into graph-level tasks as discussed in Section 3. We demonstrate the performance of node classification and link prediction on Cora and CiteSeer. For node classification, we reformulate it to graph classification by inducing an ego-graph with $10$ to $30$ nodes centered at the node to classify. Each ego-graph has the same label as the center node. For edge classification, we randomly sample 200 edges from the graph, then create 200 negative samples by replacing one node in each edge. We label each graph according to whether it is a positive or negative sample.
We use GraphCL+GCN to pre-train the GNN and report the mean performance in Table 2. The results are consistent with graph-level performance, where ImprovedAIO and Morpher outperform existing methods, with Morpher achieving slightly better performance than ImprovedAIO. Additionally, the training of the original AIO fails on both datasets due to the sparse node feature vectors.
@@ 5.3 Domain Transfer
Table 3: Domain Transfer Performance.
@@
A key problem of the graph foundation model is whether we can adapt the pre-trained models to other data domains. Here, we explore the potential of using Morpher for such adaptation. We pre-train GNNs on ENZYMES or CiteSeer datasets, then test the classification performance on MUTAG and PubMed and report the results in Table 3. We unify the pre-train feature dimension with the downstream feature dimension by padding zeros or SVD reduction. From the results, Morpher demonstrates the best transferability, followed by ImprovedAIO. Also, compared to the results on MUTAG in Table 1, all three methods have worse performances, because the GNNs are pre-trained on other datasets instead of MUTAG.
Figure 4: Results of novel class generalization (left); t-SNE embedding plots on CiteSeer, MSRC_21C (right). Train accuracy with train classes only is the accuracy of predicting the training graphs from the two training classes. Train accuracy with new test classes is the accuracy of predicting the training graphs from all three classes. Test Accuracy of zero-shot class is the accuracy of predicting the testing graphs from all three classes.
@@ 5.4 Zero-shot Classification Prototype
An advantage of adapting pre-trained GNNs to the semantic embedding space is that GNNs might be empowered to “reasoning”. Here, we conduct a novel experiment that generalizes GNN to an unseen class. Since no real-world data is available for this setting, we synthetically create three datasets, ZERO-Cora, ZERO-CiteSeer, and ZERO-PubMed, all from real-world connections. We aim to simulate a citation network with two research areas and an interdisciplinary research area in between. For each citation network, we randomly sample 120 nodes and induce their 2-hop ego-graphs, then replace the node features in 10 ego-graphs with $[1,0]$ and another 10 ego-graphs with $[0,1]$ to construct 20 training graph samples. For the remaining ego-graphs, we uniformly randomly replace the node features with $[1,0]$ and $[0,1]$ to construct 100 testing graph samples. We assign text labels of the first research area (e.g., “biology”) to the $[1,0]$ training graphs, the second research area (e.g., “informatics”) to the $[0,1]$ training graphs, and the interdisciplinary area (e.g., “bioinformatics”) to the testing graphs. Intuitively, the nodes with feature $[1,0]$ are papers in the first area, and other nodes with feature $[0,1]$ are in the second area, which makes the datasets rational.
For each dataset, using GraphCL+GCN, we pre-train GNNs on all graphs. Then, we train Morpher on the training graphs, only knowing the text labels of the two training classes. Since we do not have validation data in zero-shot learning, we report the results of each epoch in Figure 4 (left). We observe that, while Morpher quickly adapts the GNN to downstream training data, the CLIP-like framework can predict the graphs in the novel class with good accuracy (red curve). Moreover, the training samples can be classified correctly from training and novel classes. Before the training overfits, there is a period when Morpher can distinguish all the graphs from the training and novel classes with high accuracy.
Such zero-shot novel-class generalization ability validates Morpher’s alignment between graph embeddings and text embeddings. When Morpher is trained on two classes of graphs with text labels of biology and informatics, a graph-in-the-middle will be classified as text-in-the-middle: bioinformatics, even if “bioinformatics” is an unseen label. The correspondence of in-the-middle graphs and texts shows the benefit and novelty of Morpher. To the best of our knowledge, this is the first zero-shot classification prototype that generalizes GNN to unseen classes.

### v3

@@ 4 GNN Multi-modal Prompt Learning
As discussed in Sections 2 and 3, we introduce the text prompt $\mathbf{P}_{\theta}^{t}\in\mathbb{R}^{n_{t}\times d_{t}}$ with $n_{t}$ text prompt tokens and the graph prompt $\mathbf{P}_{\theta}^{g}\in\mathbb{R}^{n_{g}\times d}$ with $n_{g}$ graph prompt tokens. The graph prompting function $\psi_{g}(\cdot,\mathbf{P}_{\theta}^{g})$ modifies a given graph $\mathcal{G}$ into a manipulated graph $\mathcal{G}m=\psi_{g}(\mathcal{G},\mathbf{P}_{\theta}^{g})$.
@@
$\omega_{t}(t,\mathbf{P}_{\theta}^{t})=[\mathbf{P}_{\theta}^{t};f_{\phi}^{t}({\rm Tokenize}(t))]_{dim=0}$ (2)
@@
$\omega_{g}(\mathcal{G},\mathbf{P}_{\theta}^{g})=f_{\phi}^{g}(\mathcal{G}_{m})=f_{\phi}^{g}(\psi_{g}(\mathcal{G},\mathbf{P}_{\theta}^{g}))$ (3)
@@
$\mathbf{h}_{i}^{t}={\rm readout}(\omega_{t}(t_{i},\mathbf{P}_{\theta}^{t}))\in\mathbb{R}^{1\times d_{t}}$ (4)
$\mathbf{h}_{i}^{\mathcal{G}}={\rm readout}(\omega_{g}(\mathcal{G}_{i},\mathbf{P}_{\theta}^{g}))\in\mathbb{R}^{1\times d_{g}}$ (5)
@@
$\mathbf{z}_{norm,i}^{\mathcal{G}}=\frac{\mathbf{h}_{i}^{\mathcal{G}}}{||\mathbf{h}_{i}^{\mathcal{G}}||_{2}}=\frac{{\rm readout}(\omega_{g}(\mathcal{G}_{i},\mathbf{P}_{\theta}^{g}))}{||{\rm readout}(\omega_{g}(\mathcal{G}_{i},\mathbf{P}_{\theta}^{g}))||_{2}}$ (6)
$\mathbf{z}_{i}^{\mathcal{G}}={\rm Proj}_{\theta}(\mathbf{z}_{norm,i}^{\mathcal{G}})$ (7)
For the text embeddings, since for limited data the set $\mathcal{T}=\{t_{i}\}_{i=1}^{C}$ may contain texts that are semantically close as discussed in Appendix A.2, we extract a subspace in the text embedding space by normalizing the embedding as follows. We further normalize the text embeddings to the unit sphere.
@@
Finally, we use the in-batch similarity-based contrastive loss with temperature $\tau$ to train text prompts, graph prompts, and the projector.
@@
During inference stage, for an input graph $\mathcal{G}_{i}$ and text label candidates $\mathcal{T}=\{t_{i}\}_{i=1}^{C}$, we compute the embedding $\mathbf{z}_{i}^{\mathcal{G}}={\rm Proj}_{\theta}(\frac{{\rm readout}(\omega_{g}(\mathcal{G}_{i},\mathbf{P}_{\theta}^{g}))}{||{\rm readout}(\omega_{g}(\mathcal{G}_{i},\mathbf{P}_{\theta}^{g}))||_{2}})$ using trained $\mathbf{P}_{\theta}^{g}$ and ${\rm Proj}_{\theta}(\cdot)$. Then, we compute $\mathbf{z}^{t}_{i}$ as Equations 9 and 10. Finally, $\mathcal{G}_{i}$ will be classified to associate with text label ${\arg\max}_{1\leq i\leq C}(\mathbf{z}_{i}^{\mathcal{G}}\cdot\mathbf{z}_{i}^{t})$.
@@ 5 Experiments
IMP of ImprovedAIO (%) 3.89 $\uparrow$ 4.67 $\uparrow$ 0.90 $\uparrow$ 1.07 $\uparrow$ 1.41 $\uparrow$ 4.64 $\uparrow$ 2.29 $\uparrow$ 2.34 $\uparrow$
Further IMP of Morpher (%) 2.00 $\uparrow$ 1.72 $\uparrow$ 1.84 $\uparrow$ 5.01 $\uparrow$ 2.8 $\uparrow$ 2.24 $\uparrow$ 2.38 $\uparrow$ 4.64 $\uparrow$
Table 1: Few-shot graph classification performance (%). IMP (%): the average improvement (absolute value) compared to the best result among all the baseline methods. Best results are bolded and second-best results are underlined. We also compared with GPF-plus (Fang et al., 2023) and Gprompt (Liu et al., 2023d), in Appendix Table 14 due to space limit.
We show that both Morpher and the improved graph prompt more effectively adapt pre-trained GNNs to the specific downstream classification task. We use RoBERTa (Liu et al., 2019) as the LLM encoder for Morpher in the main experiments. We also validate the performance of Morpher with ELECTRA (Clark et al., 2020) and DistilBERT (Sanh et al., 2019) in Appendix B.3.
Datasets. We use real-world graph datasets from PyTorch Geometric (Fey and Lenssen, 2019), including one molecular dataset MUTAG (Morris et al., 2020); two bioinformatic datasets ENZYMES and PROTEINS (Borgwardt et al., 2005); one computer vision dataset MSRC_21C (Neumann et al., 2016); three citation network datasets Cora, CiteSeer and PubMed (Yang et al., 2016). We use real-world class names as text labels. The text supervision is extremely weak, as each text label contains no more than five words. More details are summarized in Appendix A.
Pre-training algorithms and GNN backbones. To pretrain GNNs for evaluation, we adopt GraphCL (You et al., 2020) and SimGRACE (Xia et al., 2022) to pre-train three widely used GNN backbones: GCN (Kipf and Welling, 2016), GAT (Yun et al., 2019) and GraphTransformer (GT) (Khosla et al., 2020). Additionally, in Appendix B.4, we verify the effectiveness of our methods on GNNs pre-trained using GraphMAE (Hou et al., 2022) and MVGRL (Hassani and Ahmadi, 2020), two other representative GNN self-supervised learning algorithms. For each dataset, to pre-train GNNs, we leverage self-supervised learning methods on all the graphs without any label information.
Baselines. We compare our methods with the following baselines: (1) training a GNN from scratch supervised by few-shot data (“supervised”); (2) fine-tuning a task head together with pre-trained GNN (“fine-tune”). We allow GNNs to be tunable for “supervised” and “fine-tune”; (3) state-of-the-art graph prompting algorithms: All-in-one (“AIO”) (Sun et al., 2023a), which is the only graph prompting algorithm that supports multiple tasks in node-level, edge-level and graph-level to the best of our knowledge; GPF-plus (Fang et al., 2023) which prompt on graph features and Gprompt (Liu et al., 2023d) which is based on subgraph similarity.
@@ 5.1 Few-shot Learning
We investigate our improved graph prompt (“ImprovedAIO”) and Multimodal prompt (“Morpher”) to adapt frozen pre-trained GNNs using few-shot data. We focus on graph-level classification here and will further investigate the few-shot learning ability at other task levels in Section 5.2. Our few-shot learning setting is more challenging than existing works (Sun et al., 2023a, 2022) as we only allow no more than 10 labeled training and validation samples for each class. The results are shown in Table 1. By observations, given the same pre-trained GNN, our ImprovedAIO outperforms all the existing baseline methods. This improvement is attributed to its design, which restricts cross-connections, ensuring stable training and optimization. Moreover, Our Morpher can achieve an absolute further accuracy improvement over the baselines across all datasets. Its superior performance, even under extremely weak text supervision, stems from its ability to dynamically adapt and align the graph and language representation spaces with prompt learning. This flexibility enables Morpher to better leverage the semantic information from weakly-supervised text labels while preserving the structural integrity of the graph embeddings, resulting in more robust and accurate predictions.
@@ 5.2 Morpher Supports Multiple-level Tasks
Table 2: Node-level, edge-level performance. Best results are bolded and second-best results are underlined.
Inherited from AIO, our ImprovedAIO and Morpher also support adaptation to downstream tasks at node-level and edge-level, because they can be reformulated into graph-level tasks. After reformulating node classification and task classification as graph classifications by inducing ego-graphs, we use GraphCL+GCN to pre-train the GNN and report the performance in Table 2. The results are consistent with graph level, where ImprovedAIO and Morpher outperform existing methods, with Morpher achieving slightly better performance than ImprovedAIO. Notably, as analyzed in Section 3, AIO is unstable during training, for example in Table 2 (Node level) and in certain cases of Table 1 (e.g., on MSRC_21C with GraphCL or with SimGRACE+GAT). The performance fluctuations observed are consistent and reflect the same underlying issue of AIO. In contrast, our proposed ImprovedAIO effectively addresses this issue.
Figure 3: Results of novel class generalization (left); t-SNE embedding plots on CiteSeer, MSRC_21C (right). Train accuracy with train classes only is the accuracy of predicting the training graphs from the two training classes. Train accuracy with new test classes is the accuracy of predicting the training graphs from all three classes. Test Accuracy of zero-shot class is the accuracy of predicting the testing graphs from all three classes. Full-resolution figures can be found in Appendix D.
@@ 5.3 Domain Transfer
We explore the potential of using Morpher for domain adaptation. From the previous experiments, we have demonstrated that our ImprovedAIO outperforms the original AIO. Therefore, in the subsequent pages, we focus on comparing Morpher with ImprovedAIO to avoid redundancy. We pre-train GNNs on ENZYMES or CiteSeer datasets, then test the classification performance on MUTAG and PubMed and report the results in Table 3. We unify the pre-train feature dimension with the downstream feature dimension by padding zeros or SVD reduction. From the results, Morpher demonstrates the best transferability, followed by ImprovedAIO.
@@
Table 3: Domain Transfer Performance. Best results are bolded and second-best results are underlined.
@@ 5.4 Zero-shot Classification Prototype
An advantage of adapting pre-trained GNNs to the semantic embedding space is that GNNs might be empowered to “reasoning” in a CLIP style. Here, we conduct a novel experiment that generalizes GNN to an unseen class. Since no real-world data is available for this setting, we synthetically create three datasets, ZERO-Cora, ZERO-CiteSeer, and ZERO-PubMed, all from real-world connections. We aim to simulate a citation network with two research areas and an interdisciplinary research area in between. For each citation network, we randomly sample 120 nodes and induce their 2-hop ego-graphs and then replace the node features in 10 ego-graphs with $[1,0]$ and another 10 ego-graphs with $[0,1]$ to construct 20 training graph samples. For the remaining ego-graphs, we uniformly randomly replace the node features with $[1,0]$ and $[0,1]$ to construct 100 testing graph samples. We assign text labels of the first research area (e.g., “biology”) to the $[1,0]$ training graphs, the second research area (e.g., “informatics”) to the $[0,1]$ training graphs, and the interdisciplinary area (e.g., “bioinformatics”) to the testing graphs. Intuitively, the nodes with feature $[1,0]$ are papers in the first area, and nodes with feature $[0,1]$ are in the second area, which makes the datasets rational.
For each dataset, using GraphCL+GCN, we pre-train GNNs on all graphs. Then, we train Morpher on the training graphs, only knowing the text labels of the two training classes. Since we do not have validation data in zero-shot learning, we report the results of each epoch in Figure 3 (left). We observe that, while Morpher quickly adapts the GNN to downstream training data, the CLIP-like framework can predict the graphs in the novel class with good accuracy (red curve). Moreover, the training samples can be classified correctly from training and novel classes. Before the training overfits, there is a period when Morpher can distinguish all the graphs from the training and novel classes with high accuracy. Such zero-shot novel-class generalization ability validates Morpher’s alignment between graph embeddings and text embeddings. When Morpher is trained on two classes of graphs with text labels of biology and informatics, a graph-in-the-middle will be classified as text-in-the-middle: bioinformatics, even if “bioinformatics” is an unseen label.